OpenStudy (anonymous):

Find the radian measure of an angle of -280 degrees

OpenStudy (whpalmer4):

To convert from degrees to radians, divide by 360 degrees and multiply by \(2\pi\) radians. \[360^\circ = 2\pi \text{ radians}\]

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

nevermind i got it thanks though! :^)

OpenStudy (whpalmer4):

What did you get for your answer?

OpenStudy (anonymous):

i got -4.884

OpenStudy (anonymous):

and sorry i didnt read your first answer right :/

OpenStudy (whpalmer4):

Yes, that's about what I get (I got a slightly larger value, -4.88692, almost certainly due to using more digits of \(\pi\) than you did)